OpenStudy (anonymous):

f(x)= (1/2)^x and x=4

OpenStudy (anonymous):

f(4)=1/16

OpenStudy (anonymous):

1/2 ^ 4 = 1^4/2^4= 1/16

OpenStudy (anonymous):

thank you :)